Problem 1.23 (Multistep) A proton is located at <3x1010, - 4x1010, - 1x1010 m Part 1 (a) What is , the vector from the origin to the location of the proton? >m By accessing this Question Assistance, you will learn while you earn points based on the Point Potential Policy set by your instructo Part 2 (b) What is the magnitude of ? the tolerance is +/-2% By accessing this Question Assistance, you will learn while you earn points based on the Point Potential Policy set by your instructo Part 3 (c) What is 7, the unit vector (vector with magnitude 1) in the direction of T? By accessing this Question Assistance, you will learn while you earn points based on the Point Potential Policy set by your instructo
Answer #1

a) The vector from the origin (0,0,0) is

  \vec{r} = < 3*10-10 , -4*10-10 , -1*10-10 > m

b) |r| = √[(3*10-10)2 + (-4*10-10)2 + (-1*10-10 )2]

= 10-10 √(9+16+1) = 5.09 *10-10 m

c) r^ = \vec{r} / |r| = <3*10-10 , -4*10-10 , -1*10-10> /  5.09 *10-10

=<0.588 , -0.785 , -0.196 >